OpenStudy (anonymous):

Find the surface area for a sphere with a radius of 10 feet. Round to the nearest whole number. 1,256 ft2 4,189 ft2 1,089 ft2 1,568 ft2 Find the volume of a sphere with a radius of 10 feet. Round to the nearest whole number. 1,257 ft3 4,187 ft3 1,089 ft3 1,568 ft3 Find the surface area for a sphere with a radius of 7 cm. Round to the nearest whole number. 307 cm2 1,436 cm2 1,020 cm2 615 cm2

OpenStudy (anonymous):

@qweqwe123123123123111

OpenStudy (anonymous):

A = 4πr² V = 4/3 πr³

OpenStudy (anonymous):

is that too the first question??

OpenStudy (anonymous):

hey my friend says your really smart qwe

OpenStudy (anonymous):

They're the answers to all 3 of your questions.

OpenStudy (anonymous):

I'm not smart. I have google. :-)

OpenStudy (anonymous):

1256 is step a for question 1

OpenStudy (anonymous):

1256 is the answer to question 1.

OpenStudy (anonymous):

thats what i thought cause i did v and i was all liek wait what!!!

OpenStudy (anonymous):

Yes...volume will be somewhat larger... :-)

OpenStudy (anonymous):

i meant 3rd questions answer is 615.44

OpenStudy (anonymous):

Yes, 615 is question 3. :-)

OpenStudy (anonymous):

and on volume does he mean 1/3?

OpenStudy (ryan123345):

Area of Sphere Formula

OpenStudy (anonymous):

No. He means 4/3 (same as 1.33333.....) v = 4/3 pi r^3

OpenStudy (ryan123345):

\[A = \frac{ 4 }{ \pi r^2 }\]

OpenStudy (ryan123345):

so for the last one your equation would be.

OpenStudy (anonymous):

is it c?

OpenStudy (ryan123345):

\[\frac{ 4 }{ \pi(49) }\]

OpenStudy (anonymous):

C? No.

OpenStudy (anonymous):

What's pi * r^3 ?

OpenStudy (anonymous):

You still here?

OpenStudy (anonymous):

3140 sorry i wasnt there

OpenStudy (anonymous):

@tafkas77

OpenStudy (anonymous):

Hi there! You mentioned me; what can I help you with? I'm not sure which question you're looking at. :)

OpenStudy (anonymous):

the middle one

OpenStudy (anonymous):

Are you still trying to figure out the answer to #3?

OpenStudy (anonymous):

.... Hello? @niall1Dawesome You there?

OpenStudy (anonymous):

sorry my computer was lagging really bad

OpenStudy (anonymous):

no number 2 i already know 1 and 3

OpenStudy (anonymous):

@tafkas77

OpenStudy (anonymous):

Okay. I'm taking a look at this right now. :)

OpenStudy (anonymous):

Eureka! I have the answer. :) Do you know the formula to calculate this?

OpenStudy (anonymous):

okay and hey you changed your icon cool

OpenStudy (anonymous):

Ha ha, yeah. :) I felt like switching it up a bit. So do you know the formula to calculate this?

OpenStudy (anonymous):

Are you still experiencing lag?

OpenStudy (anonymous):

a little

OpenStudy (anonymous):

and well isnt the formula V = 4/3 πr³

OpenStudy (anonymous):

Okay, I think you are still experiencing a little lag. No worries, I can still help you with this. Hopefully when you see it, you'll get it! :) Perfect formula! :)

OpenStudy (anonymous):

ya your post jsut showed up

OpenStudy (anonymous):

So let's plug in the values from your problem: \[v = 4/3 * \pi * 10^3\] \[v = 4/3 * \pi * 1,000\] When you plug this all in (you can do it on a calculator), you'll find that 4/3 = 1.3. so: \[v = 1.3 * 3.14 * 1,000\] And then you'll get your answer. :)

OpenStudy (anonymous):

ok i will do that

OpenStudy (anonymous):

4082

OpenStudy (anonymous):

Sorry; my mistake. You should have rounded 1.3 and input 1.4 (since 1.3333333... is irrational) and you would have gotten 4,187 :)

OpenStudy (anonymous):

oh lol

OpenStudy (anonymous):

XD Hey, everyone makes mistakes... ugh... can't say that without thinking of Hannah Montana...

OpenStudy (anonymous):

hey can you check my answer to a question like this

OpenStudy (anonymous):

yes I sure can! :)

OpenStudy (anonymous):

Find the volume for a sphere with a radius of 7 cm. Round to the nearest whole number. 307 cm2 1,436 cm2 1,020 cm2 615 cm2 im pretty sure its b but im going to redo the math

OpenStudy (anonymous):

hehe hannah montana lol

OpenStudy (anonymous):

yea i think it is b

OpenStudy (anonymous):

You are absolutely correct. :)

OpenStudy (anonymous):

yay lol thx for helping me

OpenStudy (anonymous):

You're welcome! :D

OpenStudy (anonymous):

:) everybody makes mistakes everybody has those days 1 2 3 4 lol i had too